the following equation are there
C=20+0.6yd
i=140
g=800
tr= 600
t=.33

calculate the total autonomous expenditure

calculate equilibrium level of income .calculate govt budeget surplus
how much increase in givt exoenditure is able to level up equilibrium income to 2400
if the real target output is 2640 dollars ,then by cutting the proportional income tax eate to which level will economy reacg this level
Answer #1

yd = y + tr - T = y + 600 - ty = y + 600 - 0.33y = 0.67y + 600

C = 20 + 0.6 x (0.67y + 600) = 20 + 0.402y + 360 = 380 + 0.402y

(1) Total autonomous expenditure = 380 + 140 + 800 = 1320

(2) In equilibrium, y = C + i + G

y = 380 + 0.402y + 140 + 800

(1 - 0.402)y = 1320

0.598y = 1320

y = 2207

(3) T = 0.33 x 2207 = 728.43

Budget surplus = T - G = 728.43 - 800 = -71.57 (Indicates budget deficit)

(4) When y = 2400,

2400 = 380 + 0.402y + 140 + G

2400 = 520 + (0.402 x 2400) + G

1880 = 964.8 + G

G = 835.2

Increase in G = 835.2 - 800 = 35.2

NOTE: As per Answering Policy, 1st 4 parts are answered.